OpenStudy (anonymous):

Simplify help please I cant believe I forgot how 2 do this -14x^3 ______________ x^3 - 5x^4

OpenStudy (anonymous):

\[\frac{-14x^3}{x^3-5x^4}\] \[=\frac{-14x^3}{x^3(1-5x)}\] then you can cancel the \(x^3\)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

I cant give medals on my kindle but ill use a computer tomorrow to get one 2 u